# CAME-AB: Cross-Modality Attention with Mixture-of-Experts for Antibody Binding Site Prediction

### v1

@@ A.4 PyBioMed: Physicochemical Properties
PyBioMed is a Python-based toolkit that extracts a wide range of physicochemical and structural features from protein sequences and structures, including hydrophobicity, charge, and secondary structure propensity ([22]). These descriptors provide a rich feature set for downstream tasks, complementing sequence and evolutionary representations.
@@ A.5 Multi-view Learning in Bioinformatics
Multi-view learning integrates complementary information from diverse feature representations, enabling more robust and accurate predictions in bioinformatics. For example, combining amino acid sequence features ([6]), evolutionary profiles (e.g., BLOSUM and ESMC), and physicochemical properties ([37]) has consistently demonstrated superior performance in tasks such as protein function prediction and binding site identification. Recent advances in machine learning models, including Graph Neural Networks (GNNs) and Transformer-based architectures, have further enhanced multi-view learning by effectively capturing the relationships between different feature views.
@@ Appendix B Evaluation Metrics
    where $TP$ and $FP$ denote the number of true positive and false positive predictions, respectively.
@@
    where $FN$ is the number of false negatives.
@@
    where $TN$ denotes true negatives.

### v2

@@ A.4 PyBioMed: Physicochemical Properties
PyBioMed is a Python-based toolkit that extracts a wide range of physicochemical and structural features from protein sequences and structures, including hydrophobicity, charge, and secondary structure propensity [22]. These descriptors provide a rich feature set for downstream tasks, complementing sequence and evolutionary representations.
@@ A.5 Multi-view Learning in Bioinformatics
Multi-view learning integrates complementary information from diverse feature representations, enabling more robust and accurate predictions in bioinformatics. For example, combining amino acid sequence features [6], evolutionary profiles (e.g., BLOSUM and ESMC), and physicochemical properties [37] has consistently demonstrated superior performance in tasks such as protein function prediction and binding site identification. Recent advances in machine learning models, including Graph Neural Networks (GNNs) and Transformer-based architectures, have further enhanced multi-view learning by effectively capturing the relationships between different feature views.
@@ Appendix B Evaluation Metrics
    where TPTP and FPFP denote the number of true positive and false positive predictions, respectively.
@@
    where FNFN is the number of false negatives.
@@
    where TNTN denotes true negatives.